# Inner non-Hermitian skin effect on Bethe lattice

### v1

@@ Abstract
Introduction.- The Bethe lattice is a tree-like geometry, in which each vertex has the same number of neighbors. Its key feature is the absence of closed loops, which reduces complex quantum interference effects and makes lattice models on the Bethe lattice generally easier to solve compared to other latticesPhysRevB.63.155110; Aryal_2020. For instance, simple models like the Ising model can be solved exactly on the Bethe latticeda Silva and Coutinho (1986). Additionally, the study of interesting models on the Bethe lattice is motivated by the similarity of results to those on crystal lattices. Solutions on the Bethe lattice often provide a better approximation than the mean-field approach for thermodynamic quantities of models on regular latticesGujrati (1995); Eckstein et al. (2005), such as square and honeycomb lattices, offering valuable insights into the physical properties of lattice models. Consequently, significant advances have been made in studying various models on the Bethe lattice, including classical statistical models, quantum spin models, spin-glass systems da Silva and Coutinho (1986); Thouless (1986); Tragtenberg and Yokoi (1995); Laumann et al. (2008); Semerjian et al. (2009); Laumann et al. (2009); Li et al. (2012); Daniška and Gendiar (2016); Qu et al. (2019); Lunts et al. (2021), the Bose/Fermi Hubbard model, and the Anderson model Zirnbauer (1986); Mirlin and Fyodorov (1994); De Luca et al. (2014); Savitz et al. (2019).
@@
In recent years, there has been a remarkable surge in research interest in non-Hermitian physicsYao and Wang (2018); Lieu (2018); PhysRevLett.121.136802; GongZ18prx; PhysRevLett.123.246801; LiuT19prl; JinL19prb; Kawabata19prx; PhysRevLett.123.090603; Longhi (2019); Yuce and Ramezani (2019); Chen et al. (2019); Yu et al. (2020b); Henriques et al. (2020); Zeng20prb; HuH21prl; GuoCX21prl; TianyuLi2021; Bergholtz21rmp; reviewskineffect; PhysRevB.105.094103; PhysRevB.108.245114; PhysRevLett.133.076502; PRXwangzhong; WanerHou2024, owing to its rich phenomena and promising applications. As the simplest non-Hermitian topological system, the non-Hermitian version of the SSH model has been widely investigatedLieu (2018); Yuce and Ramezani (2019); Henriques et al. (2020); Chen et al. (2019); Yu et al. (2020b), revealing many intriguing properties such as the non-Hermitian skin effect and the breakdown of conventional bulk-boundary correspondenceYao and Wang (2018); PhysRevLett.123.246801. To address these issues, non-Bloch band theory has been constructed in one dimension based on the generalized Brillouin zone, and it has been successfully applied to various non-Hermitian systemsPhysRevLett.123.066404; PhysRevLett.121.136802. In two and higher dimensions, the skin effect can be manifested in different forms, such as the corner-skin effect and geometry-dependent skin effectLonghi (2019); Lee et al. (2019); Kawabata et al. (2020); Okugawa et al. (2020); FuY21prb; Zhangx21nc; Li et al. (2022); Zhu and Gong (2022); Li et al. (2023); Sun et al. (2023). It is proposed that the existence of the skin effect depends on whether the periodic-boundary spectrum of the Hamiltonian covers a finite area in the complex planeZhang2022. For system of arbitrary spatial dimensions, an amoeba theory for the non-Hermitian skin effect and non-Bloch band theory has also been developedPRXwangzhong. For aforementioned Bethe lattice with fractal structures, we have recently shown that the non-Hermiticity can give rise to particular quantum fractals in terms of energy and wavefunctions on itSun24NHQF, while the application of more general non-Hermitian effects and possible skin effects on Bethe lattice are largely unexplored.
@@
In this work, we implement the non-Hermitian SSH model on Bethe lattice and explore its intriguing non-Hermitian features which arise from the interplay between non-Hermitian topology and the self-similar lattice structure. We analytically solve the non-Hermitian SSH model on Bethe lattice by developing a decoupling technique, obtaining the energy spectrum and corresponding eigenstates. We identify a distinct form of non-Hermitian skin effect with eigenstates localizing within the bulk region, coined ‘inner non-Hermitian skin effect’innerskin. We explain this particular inner non-Hermtian skin effect from the combination of non-Hermiticity and bulk-boundary duality of Bethe lattice. To further show the experimental relevance of our proposal, we implement the non-Hermitian SSH model using a Bethe-lattice circuit, which further validates our theoretical findings.
@@
Inner non-Hermitian skin effect.- The skin effect is an unusual phenomenon in non-Hermitian systems, where all eigenstates tend to localize near boundaries of the system under open boundary conditions. Interestingly, on the Bethe lattice, a different form of non-Hermitian skin effect arises, with eigenstates localizing within the bulk region instead. Explicitly, as illustrated in Fig. 2 (d), various categories of eigenstates localize near different generations within the central region. We refer to such anomalous phenomenon as inner non-Hermitian skin effectinnerskin. This behavior shows stark difference from the non-Hermitian skin effect on regular lattices, in which all of the skin modes localize at boundaries. Note that the topologically protected edge states from SSH model emerging at zero energy exhibit similar localization behaviors, but with a higher degree of localization.
@@
When $|t_{1}|>|\gamma|$, the 1D non-Hermitian SSH model can be transformed into a Hermitian one, with the alternating hopping amplitudes being $\bar{t}_{1},\bar{t}_{2}$. For a general SSH model given byjps2017:
@@
Simulations of inner skin effect on electric circuit.- To further validate our proposal, we implement the non-Hermitian SSH model using a Bethe-lattice circuit composed of capacitors, inductors, and operational amplifiers. The non-Hermitian effects resulting from non-reciprocal hopping can be incorporated by utilizing operational amplifiersLee2018; PhysRevResearch.3.023056; PhysRevB.103.125411; PhysRevB.103.035420; Liu2022; PhysRevResearch.5.013107. By applying Kirchhoff’s current law, we can design a circuit that corresponds to our non-Hermitian tight-binding model in Eq. (Inner non-Hermitian skin effect on Bethe lattice), with the basic circuit elements depicted in Fig. 3. The Laplacian operator $\mathcal{L}$ for this circuit, which relates the current and voltage distributions in the circuit ($I=\mathcal{L}V$), can be written as:
@@
Note added.- After completing the present manuscript, we came to notice related works on arXiv:2408.11024multifractal and arXiv:2409.01873quantumtransportbethe.

### v2

@@ Abstract
Introduction.- The Bethe lattice is a tree-like geometry, in which each vertex has the same number of neighbors. Its key feature is the absence of closed loops, which reduces complex quantum interference effects and makes lattice models on the Bethe lattice generally easier to solve compared to other latticesMahan (2001); Aryal and Kettemann (2020). For instance, simple models like the Ising model can be solved exactly on the Bethe latticeda Silva and Coutinho (1986). Additionally, the study of interesting models on the Bethe lattice is motivated by the similarity of results to those on crystal lattices. Solutions on the Bethe lattice often provide a better approximation than the mean-field approach for thermodynamic quantities of models on regular latticesGujrati (1995); Eckstein et al. (2005), such as square and honeycomb lattices, offering valuable insights into the physical properties of lattice models. Consequently, significant advances have been made in studying various models on the Bethe lattice, including classical statistical models, quantum spin models, spin-glass systems da Silva and Coutinho (1986); Thouless (1986); Tragtenberg and Yokoi (1995); Laumann et al. (2008); Semerjian et al. (2009); Laumann et al. (2009); Li et al. (2012); Daniška and Gendiar (2016); Qu et al. (2019); Lunts et al. (2021), the Bose/Fermi Hubbard model, and the Anderson model Zirnbauer (1986); Mirlin and Fyodorov (1994); De Luca et al. (2014); Savitz et al. (2019).
@@
In recent years, there has been a remarkable surge in research interest in non-Hermitian physicsYao and Wang (2018); Lieu (2018); Yao et al. (2018); Gong et al. (2018); Song et al. (2019); Liu et al. (2019); Jin and Song (2019); Kawabata et al. (2019); Hamazaki et al. (2019); Longhi (2019); Yuce and Ramezani (2019); Chen et al. (2019); Yu et al. (2020); Henriques et al. (2020); Zeng et al. (2020); Hu and Zhao (2021); Guo et al. (2021); Li et al. (2021); Bergholtz et al. (2021); Xiujuan Zhang and Chen (2022); Cheng et al. (2022); Hu et al. (2023); Yoshida et al. (2024); Wang et al. (2024); Hou et al. (2024), owing to its rich phenomena and promising applications. As the simplest non-Hermitian topological system, the non-Hermitian version of the SSH model has been widely investigatedLieu (2018); Yuce and Ramezani (2019); Henriques et al. (2020); Chen et al. (2019); Yu et al. (2020), revealing many intriguing properties such as the non-Hermitian skin effect and the breakdown of conventional bulk-boundary correspondenceYao and Wang (2018); Song et al. (2019). To address these issues, non-Bloch band theory has been constructed in one dimension based on the generalized Brillouin zone, and it has been successfully applied to various non-Hermitian systemsYokomizo and Murakami (2019); Yao et al. (2018). In two and higher dimensions, the skin effect can be manifested in different forms, such as the corner-skin effect and geometry-dependent skin effectLonghi (2019); Lee et al. (2019); Kawabata et al. (2020); Okugawa et al. (2020); Fu et al. (2021); Zhang et al. (2021); Li et al. (2022); Zhu and Gong (2022); Li et al. (2023); Sun et al. (2023). It is proposed that the existence of the skin effect depends on whether the periodic-boundary spectrum of the Hamiltonian covers a finite area in the complex planeZhang et al. (2022). For system of arbitrary spatial dimensions, an amoeba theory for the non-Hermitian skin effect and non-Bloch band theory has also been developedWang et al. (2024). For aforementioned Bethe lattice with fractal structures, we have recently shown that the non-Hermiticity can give rise to particular quantum fractals in terms of energy and wavefunctions on itSun et al. (2024), while the application of more general non-Hermitian effects and possible skin effects on Bethe lattice are largely unexplored.
@@
In this work, we implement the non-Hermitian SSH model on Bethe lattice and explore its intriguing non-Hermitian features which arise from the interplay between non-Hermitian topology and the self-similar lattice structure. We analytically solve the non-Hermitian SSH model on Bethe lattice by developing a decoupling technique, obtaining the energy spectrum and corresponding eigenstates. We identify a distinct form of non-Hermitian skin effect with eigenstates localizing within the bulk region, coined ‘inner non-Hermitian skin effect’Manna and Roy (2023). We explain this particular inner non-Hermtian skin effect from the combination of non-Hermiticity and bulk-boundary duality of Bethe lattice. To further show the experimental relevance of our proposal, we implement the non-Hermitian SSH model using a Bethe-lattice circuit, which further validates our theoretical findings.
@@
Inner non-Hermitian skin effect.- The skin effect is an unusual phenomenon in non-Hermitian systems, where all eigenstates tend to localize near boundaries of the system under open boundary conditions. Interestingly, on the Bethe lattice, a different form of non-Hermitian skin effect arises, with eigenstates localizing within the bulk region instead. Explicitly, as illustrated in Fig. 2 (d), various categories of eigenstates localize near different generations within the central region. We refer to such anomalous phenomenon as inner non-Hermitian skin effectManna and Roy (2023). This behavior shows stark difference from the non-Hermitian skin effect on regular lattices, in which all of the skin modes localize at boundaries. Note that the topologically protected edge states from SSH model emerging at zero energy exhibit similar localization behaviors, but with a higher degree of localization.
@@
When $|t_{1}|>|\gamma|$, the 1D non-Hermitian SSH model can be transformed into a Hermitian one, with the alternating hopping amplitudes being $\bar{t}_{1},\bar{t}_{2}$. For a general SSH model given byEliashvili et al. (2017):
@@
Simulations of inner skin effect on electric circuit.- To further validate our proposal, we implement the non-Hermitian SSH model using a Bethe-lattice circuit composed of capacitors, inductors, and operational amplifiers. The non-Hermitian effects resulting from non-reciprocal hopping can be incorporated by utilizing operational amplifiersLee et al. (2018); Dong et al. (2021); Xu et al. (2021); Rafi-Ul-Islam et al. (2021); Liu et al. (2022); Rafi-Ul-Islam et al. (2023). By applying Kirchhoff’s current law, we can design a circuit that corresponds to our non-Hermitian tight-binding model in Eq. (Inner non-Hermitian skin effect on Bethe lattice), with the basic circuit elements depicted in Fig. 3. The Laplacian operator $\mathcal{L}$ for this circuit, which relates the current and voltage distributions in the circuit ($I=\mathcal{L}V$), can be written as:
@@
Note added.- After completing the present manuscript, we came to notice related works on arXiv:2408.11024Hamanaka et al. (2024) and arXiv:2409.01873Hatano et al. (2024).